OpenStudy (anonymous):

The average of 5 consecutive even numbers A,B,C,D and E is 52. What is the product of B and E?

OpenStudy (anonymous):

implies that c=52 b is then 50 and e is 56 50*56?

OpenStudy (anonymous):

answer plz

OpenStudy (anonymous):

2800

OpenStudy (anonymous):

x x+2 x+4 like that

OpenStudy (anonymous):

\[ \frac{a + (a + 2) +(a + 4) + (a + 6) + 52 }{ 5 }\]

OpenStudy (anonymous):

why c=52

OpenStudy (anonymous):

since it is the average of consecutive terms, the middle term is just the average

OpenStudy (anonymous):

for every question like that ya

OpenStudy (anonymous):

next is a+8 la u put 52

OpenStudy (anonymous):

oh im supposing x as c, not a

OpenStudy (anonymous):

\[ \frac{a+(a+2)+(a+4)+(a+6)+52}{5} \] A = A B = (a + 2) C = (a +4) D = (a +6) E = (a + 8) = 52 E = 52 D = 44 = (a = 52 - 8) C = 38 = (a = 44 - 6) B = 34 = (a = 38 - 4) A = 32 = (a = 34 - 2) Product of B & E = 34 * 52 = 1768

OpenStudy (anonymous):

wtf is thsi^

OpenStudy (anonymous):

ohhh wait its the average of the 5 numbers is 52 or e is 52?

OpenStudy (anonymous):

It could be both ways. lol anyways, she has the solution for both lol

OpenStudy (anonymous):

If the average of a+b+c+d+e = 52 I would have worded it like The average of 5 consecutive even numbers A,B,C,D,E is 52 Anyways, it could be read both ways but the solution to both are here just in case anyone else looks it up and needs help